OpenStudy (anonymous):

A tuition scale for a state university shows an increase of $1,100 each year. Part 1: If the initial cost was $27,000, what will the cost of tuition be in twenty two years? (5 points) Part 2: In complete sentences, explain how an arithmetic sequence was applied. (4 points)

OpenStudy (anonymous):

1: 27000 + 1100n where n=22

OpenStudy (tkhunny):

"$1100 each year" These words should cause alarms to go off in your head. This is an arithmetic progression (sequence)! The change is the same each year. 1 year 1100 (1100*1) 2 years 2200 (1100*2) 3 years 3300 (1100*3) and so forth,

OpenStudy (anonymous):

is that right ?

OpenStudy (tkhunny):

Start with 27000 instead of 2700 and you'll have it.

OpenStudy (anonymous):

27,000+1100n n=22 27,00+24200 26900

OpenStudy (tkhunny):

Nope. 2700 reappeared 27000 + 24200 = ??

OpenStudy (anonymous):

51200

OpenStudy (tkhunny):

That's it. Do you not like the number for some reason?

OpenStudy (anonymous):

lol idk :(. i have one more . do you mind helping ?

OpenStudy (tkhunny):

Did we answer part 2?

OpenStudy (anonymous):

ohh no i didnt :(

OpenStudy (tkhunny):

Ha! I actually stuck it in up front. Go read my first post. See if you find the clue.

OpenStudy (anonymous):

use the progression sequence ?

OpenStudy (tkhunny):

Next part. The change is constant. The first change is some value. The second change is the same value. The 33rd change is the same value. Every year the change is the same!